OpenStudy (anonymous):

evaluate the integral of x e^(-x^2) from minus infinity to infinity. I already found the integral but I don't understand why the limits of this will give 0

OpenStudy (amistre64):

well, e^(-x^2) is an even function and x is an odd function even times odd = odd

OpenStudy (amistre64):

the integral from 0 to infinity is the opposite of it from 0 to -inf

OpenStudy (anonymous):

@cpham did you try drawing it out... ? The key is the symmetry about \(y=x\)... observe:$$f(x)=xe^{-x^2}\\f(-x)=(-x)e^{-(-x)^2}=-xe^{-x^2}=-f(x)$$